find the variance and standard deviation of the probability distribution of the random variable W if P{W = w} = {w+1}/20 for W = {1, 2, 3, 4, 5}


1
Expert's answer
2022-03-22T04:23:41-0400

"V(W)=M(W^2)-M^2(W)"

"P(W=1)={\frac 2 {20}}"

"P(W=2)={\frac 3 {20}}"

"P(W=3)={\frac 4 {20}}"

"P(W=4)={\frac 5 {20}}"

"P(W=5)={\frac 6 {20}}"

"M(W)={\frac 2 {20}}*1+{\frac 3 {20}}*2+{\frac 4 {20}}*3+{\frac 5 {20}}*4+{\frac 6 {20}}*5={\frac {70}{20}}=3.5\implies M^2(W)=12.25"

"M(W^2)={\frac 2 {20}}*1^2+{\frac 3 {20}}*2^2+{\frac 4 {20}}*3^2+{\frac 5 {20}}*4^2+{\frac 6 {20}}*5^2={\frac {280}{20}}=14"

So, "V(W)=14-12.25=1.75"

"\sigma(W)=\sqrt{V(W)}=\sqrt{1.75}\approx1.323"